# Radiative Cooling Effects on Plasmoid Formation in Black Hole Accretion Flows with Multiple Magnetic Loops

### v1

@@ 3.2.5 Energy Extraction
Plasmoid chain created by magnetic reconnection in BH ergosphere is one of the possible mechanisms of energy extraction from a rotating BH via the Penrose process (PP) (e.g., Asenjo and Comisso 2015). In previous studies of reconnection-driven PP, the discussion has primarily focused on current sheets forming near the equatorial plane of the BH (Koide and Arai 2008; Asenjo and Comisso 2015). A recent study of Camilloni and Rezzolla (2025) has shown that PP can also occur outside the equatorial plane, as strong magnetic reconnection regions are present on the transition region between the accretion disk and the jet (Dimitropoulos et al. 2025). To precisely characterize PP, a local model of magnetic reconnection is required, providing a prescription for the plasmoid outflow velocity field $\tilde{v}_{\text{out}}$ and their associated energetics. However, the quantities of most importance, outflow velocity and the reconnection rate, come from MHD approximation (Shen and YuChih 2025; Camilloni and Rezzolla 2025). In our study, we do not pursue further theoretical analysis. Instead, we focus on how radiative cooling impact the energy extraction process via magnetic reconnection although it happens numerically due to numerical resistivity in ideal GRMHD simulations.
@@
To present directly that the plasmoid carrying negative energy-at-infinity actually falls into the BH, we track the evolution of a selected plasmoid near the equatorial plane, as shown in Fig. 15. The time evolution clearly shows the inward motion of the plasmoid with negative energy-at-infinity as it plunges into the BH.
As discussed in Sect. 3.2.3, radiative cooling affects the local reconnection rate and the efficiency of plasmoid formation. As shown in Fig. 16, we adopt previously introduced $hu_{t}$ as the diagnostic criterion to identify negative-energy regions within plasmoids, and compute the total negative energy by summing over these regions for each model. In model A (strong radiative cooling), two prominent peaks are observed around $t=14050\,\rm M$ and $t=14700\,\rm M$. Similarly, in Fig. 10, peaks corresponding to plasmoid formation events are observed near these times, further supporting the temporal correlation. For Model C (no cooling), although no prominent peaks are observed, there still exist noticeable variations corresponding to plasmoid formation events between $t=14200\,\rm M$ and $t=14600\,\rm M$, compared to the surrounding periods. Model A exhibits a higher average negative energy-at-infinity density compared to the non-cooling model C. Thus, to some extent, radiative cooling may lead to more efficient energy extraction. Although a rigorous calculation of energy extraction requires a more detailed discussion, the trend-based analysis of negative energy differences across different models presented here provide supporting evidence for our discussion on the effects of radiative cooling.
@@ 4 Summary and Discussion
    Radiative cooling results in a higher frequency of plasmoid formation. However, the average “lifetime” of individual plasmoids decreases due to cooling-induced collapse. Radiative cooling case tends to have more small-scale reconnection events and radiative plasmoid collapses. The interplay between enhanced reconnection and rapid collapse maintains a similar average number of plasmoids.
@@ Acknowledgements.
To investigate the effects of numerical resolution, we conduct two-dimensional GRMHD simulations of magnetized accretion flows onto a rotating BH, incorporating radiative cooling, at different grid resolutions. Model $\tt{hr}$ has the same initial condition with model A, only the resolution is different. We adopt a grid size of $(N_{r},N_{\theta})=1024\times 512$ as the standard resolution and We adopted five levels of static mesh refinement with a base resolution of $(N_{r},N_{\theta})=256\times 128$. The second refinement level is applied within the radial range $100<r/r_{g}<200$, the third within $90<r/r_{g}<100$, the fourth within $60<r/r_{g}<90$, and the fifth within $3<r/r_{g}<60$. As a result, the effective resolution in the regions of prime interest reaches $(N_{r},N_{\theta})=4096\times 2048$, allowing us to capture the dynamics near the current sheet with significantly enhanced spatial resolution. As shown in Fig. B17, the evolutionary trends between different resolution cases show excellent agreement.

### v2

@@ 3.2.5 Energy Extraction
Plasmoid chain created by magnetic reconnection in BH ergosphere is one of the possible mechanisms of energy extraction from a rotating BH via the Penrose process (PP) (e.g., Asenjo and Comisso 2015). In previous studies of reconnection-driven PP, the discussion has primarily focused on current sheets forming near the equatorial plane of the BH (Koide and Arai 2008; Asenjo and Comisso 2015). A recent study of Camilloni and Rezzolla (2025) has shown that PP can also occur outside the equatorial plane, as strong magnetic reconnection regions are present on the transition region between the accretion disk and the jet (Dimitropoulos et al. 2025). To precisely characterize PP, a local model of magnetic reconnection is required, providing a prescription for the plasmoid outflow velocity field $\tilde{v}_{\text{out}}$ and their associated energetics. However, the quantities of greatest importance, outflow velocity and the reconnection rate, come from MHD approximation (Shen and YuChih 2025; Camilloni and Rezzolla 2025). In our study, we do not pursue further theoretical analysis. Instead, we focus on how radiative cooling impact the energy extraction process via magnetic reconnection although it happens numerically due to numerical resistivity in ideal GRMHD simulations.
@@
To present directly that the plasmoid carrying negative energy-at-infinity falls into the BH, we track the evolution of a selected plasmoid near the equatorial plane, as shown in Fig. 15. The time evolution clearly shows the inward motion of the plasmoid with negative energy-at-infinity as it plunges into the BH.
As discussed in Sect. 3.2.3, radiative cooling affects the local reconnection rate and the efficiency of plasmoid formation. As shown in Fig. 16, we adopt previously introduced $hu_{t}$ as the diagnostic criterion to identify negative-energy regions within plasmoids, and compute the total negative energy by summing over these regions for each model. In model A (strong radiative cooling), two prominent peaks are observed around $t=14050\,\rm M$ and $t=14700\,\rm M$. Similarly, in Fig. 10, peaks corresponding to plasmoid formation events are observed near these times, further supporting the temporal correlation. For Model C (no cooling), although no prominent peaks are observed, there still exists noticeable variations corresponding to plasmoid formation events between $t=14200\,\rm M$ and $t=14600\,\rm M$, compared to the surrounding periods. Model A exhibits a higher average negative energy-at-infinity density compared to the non-cooling model C. Thus, to some extent, radiative cooling may lead to more efficient energy extraction. Although a rigorous calculation of energy extraction requires a more detailed discussion, the trend-based analysis of negative energy differences across different models presented here provide supporting evidence for our discussion on the effects of radiative cooling.
@@ 4 Summary and discussion
    Radiative cooling results in a higher frequency of plasmoid formation. However, the average “lifetime” of individual plasmoids decreases due to cooling-induced collapse. The radiative cooling case tends to have more small-scale reconnection events and radiative plasmoid collapses. The interplay between enhanced reconnection and rapid collapse maintains a similar average number of plasmoids.
@@ Appendix B Resolution test
To investigate the effects of numerical resolution, we conduct two-dimensional GRMHD simulations of magnetized accretion flows onto a rotating BH, incorporating radiative cooling, at different grid resolutions. Model $\tt{hr}$ has the same initial condition as model A, differing only in resolution. We adopt a grid size of $(N_{r},N_{\theta})=1024\times 512$ as the standard resolution and We adopted five levels of static mesh refinement with a base resolution of $(N_{r},N_{\theta})=256\times 128$. The second refinement level is applied within the radial range $100<r/r_{g}<200$, the third within $90<r/r_{g}<100$, the fourth within $60<r/r_{g}<90$, and the fifth within $3<r/r_{g}<60$. As a result, the effective resolution in the regions of prime interest reaches $(N_{r},N_{\theta})=4096\times 2048$, allowing us to capture the dynamics near the current sheet with significantly enhanced spatial resolution. As shown in Fig. B17, the evolutionary trends between different resolution cases show excellent agreement.